# An Adaptive Online Smoother with Closed-Form Solutions and Information-Theoretic Lag Selection for Conditional Gaussian Nonlinear Systems

### v1

@@ 2.2 Analytically Solvable Filter and Smoother Posterior Distributions
Then the system of optimal nonlinear smoother equations is given by the following backward differential equations:
$\overleftarrow{{\,\rm d}\boldsymbol{\mu}_{\text{{s}}}}$ $=-\left(\boldsymbol{\Lambda}^{\mathbf{y}}\boldsymbol{\mu}_{\text{{s}}}+\mathbf{f}^{\mathbf{y}}-\mathbf{B}\mathbf{R}_{\text{{f}}}^{-1}(\boldsymbol{\mu}_{\text{{f}}}-\boldsymbol{\mu}_{\text{{s}}})\right){\,\rm d}t+(\boldsymbol{\Sigma}^{\mathbf{y}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})(\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})^{-1}\left(\overleftarrow{{\,\rm d}\mathbf{x}}+(\boldsymbol{\Lambda}^{\mathbf{x}}\boldsymbol{\mu}_{\text{{s}}}+\mathbf{f}^{\mathbf{x}}){\,\rm d}t\right),$ (2.8)
$\overleftarrow{{\,\rm d}\mathbf{R}_{\text{{s}}}}$ $=-\left((\mathbf{A}+\mathbf{B}\mathbf{R}_{\text{{f}}}^{-1})\mathbf{R}_{\text{{s}}}+\mathbf{R}_{\text{{s}}}(\mathbf{A}+\mathbf{B}\mathbf{R}_{\text{{f}}}^{-1})^{\dagger}-\mathbf{B}\right){\,\rm d}t,$ (2.9)
for $T\geq t\geq 0$, with $t$ running backward. The backward-arrow notation on the left-hand side of (2.8)–(2.9) and in $\overleftarrow{{\,\rm d}\mathbf{x}}$ on right-hand side of (2.8) is to be understood as
$\overleftarrow{{\,\rm d}\boldsymbol{\mu}_{\text{{s}}}}:=\lim_{\Delta t\to 0}\left(\boldsymbol{\mu}_{\text{{s}}}(t)-\boldsymbol{\mu}_{\text{{s}}}(t+\Delta t)\right),\quad\overleftarrow{{\,\rm d}\mathbf{R}_{\text{{s}}}}:=\lim_{\Delta t\to 0}\left(\mathbf{R}_{\text{{s}}}(t)-\mathbf{R}_{\text{{s}}}(t+\Delta t)\right),\quad\overleftarrow{{\,\rm d}\mathbf{x}}:=\lim_{\Delta t\to 0}\left(\mathbf{x}(t)-\mathbf{x}(t+\Delta t)\right),$
for $\boldsymbol{\mu}_{\text{{s}}}(t)$ and $\mathbf{R}_{\text{{s}}}(t)$ being the conditional mean and covariance of the smoother posterior distribution at time $t$, which is Gaussian,
$\mathbb{P}\big{(}\mathbf{y}(t)\big{|}\mathbf{x}(s),s\in[0,T]\big{)}\overset{{\,\rm d}}{\sim}\mathcal{N}_{l}(\boldsymbol{\mu}_{\text{{s}}}(t),\mathbf{R}_{\text{{s}}}(t)).$
In other words, the notation $\overleftarrow{\frac{{\,\rm d}\boldsymbol{\cdot}}{{\,\rm d}t}}$ corresponds to the negative of the usual derivative, which means the system in (2.8)–(2.9) is meant to be solved backward over $[0,T]$. The “starting” values for the smoother posterior statistics, $\left(\boldsymbol{\mu}_{\text{{s}}}(T),\mathbf{R}_{\text{{s}}}(T)\right)$, are the same as those of the corresponding filter estimates at the endpoint $t=T$, $\left(\boldsymbol{\mu}_{\text{{f}}}(T),\mathbf{R}_{\text{{f}}}(T)\right)$.
@@
This is Theorem 12.10 in Liptser & Shiryaev [36]. For the analogous result in the case of discrete time, see Theorem 13.12, which outlines the corresponding optimal recursive nonlinear smoother backward difference equations, with the respective sufficient assumptions given in Subchapters 13.2.1 and 13.3.8. For a martingale-free proof see Andreou & Chen [68] (although it requires the additional assumptions (10)(12) found in Appendix A, as to pass from the discrete-time smoother to the continuous one via a formal limit). ∎
@@ 3 Optimal Online Smoother
The standard smoothing procedure involves a forward pass using a filtering method followed by a backward pass to obtain the optimal smoother state estimation [15, 16, 17, 18, 19]. While this forward-backward algorithm is effective for offline data postprocessing, in many practical situations, new observations become available sequentially. Recalculating the smoother’s posterior Gaussian statistics from scratch each time is computationally and storage-intensive. Therefore, an online version of the optimal nonlinear smoother, requiring only forward-in-time model updates, is highly desirable, as it can update estimated states as new data arrive. While particle filters and ensemble methods offer approximate state estimates for general CTNDSs, the CGNS framework provides exact, closed-form analytic solutions for deriving an online smoother. The online optimal nonlinear smoother not only efficiently updates the estimates but also facilitates the understanding of the mathematical and numerical properties of the system.
@@ 3.1 Optimal Online Forward-In-Time Discrete Smoother
In what follows, the superscript notation $\boldsymbol{\cdot}^{\,j}$ is used to denote the discrete approximation to the continuous form of the respective vector or matrix functional when evaluated on $t_{j}$, for example $\boldsymbol{\Lambda}^{\mathbf{x},j}:=\boldsymbol{\Lambda}^{\mathbf{x}}(t_{j},\mathbf{x}(t_{j}))$. We write the CGNS of equations (2.1)–(2.2) in a discrete fashion using the Euler-Maruyama scheme [74, 75, 76]. Then the time discretization of (2.1)–(2.2) is simply given by
@@
$\boldsymbol{\mu}_{\text{{s}}}^{j,n}:=\mathbb{E}\left[\mathbf{y}^{j}\big{|}\mathbf{x}^{s},s\leq n\right],\quad\mathbf{R}_{\text{{s}}}^{j,n}:=\mathrm{Cov}\left(\mathbf{y}^{j},\mathbf{y}^{j}\big{|}\mathbf{x}^{s},s\leq n\right)=\mathbb{E}\left[(\mathbf{y}^{j}-\boldsymbol{\mu}_{\text{{s}}}^{j,n})(\mathbf{y}^{j}-\boldsymbol{\mu}_{\text{{s}}}^{j,n})^{\dagger}\big{|}\mathbf{x}^{s},s\leq n\right].$
@@
Notice how we explicitly note the dependence of the smoother state estimates on the length of the observational period. As is known by the discrete-time counterpart of Theorem 2.1 (see Theorem 13.3 in Liptser & Shiryaev [36]), the smoother posterior distribution $\mathbb{P}\big{(}\mathbf{y}^{j}\big{|}\mathbf{x}^{s},s=0,\ldots,n\big{)}$ is ($\mathbb{P}$-a.s.) conditionally Gaussian, $\mathcal{N}_{l}(\boldsymbol{\mu}_{\text{{s}}}^{j,n},\mathbf{R}_{\text{{s}}}^{j,n})$, and it is possible to show that the conditional mean $\boldsymbol{\mu}_{\text{{s}}}^{j,n}$ and conditional covariance $\mathbf{R}_{\text{{s}}}^{j,n}$ of the discrete smoother at time step $t_{j}$ when conditioning up to the $n$-th observation satisfy the following recursive backward difference equations for $n\in\mathbb{N}$ and $j=0,1,\ldots,n-1$ [68] (under regularity assumptions (1)(12) in Appendix A and all other assumed conditions):
$\boldsymbol{\mu}_{\text{{s}}}^{j,n}$ $=\boldsymbol{\mu}_{\text{{f}}}^{j}+\mathbf{E}^{j}\left(\boldsymbol{\mu}_{\text{{s}}}^{j+1,n}-(\mathbf{I}_{l\times l}+\boldsymbol{\Lambda}^{\mathbf{y},j}\Delta t)\boldsymbol{\mu}_{\text{{f}}}^{j}-\mathbf{f}^{\mathbf{y},j}\Delta t\right)+\mathbf{F}^{j}\left(\mathbf{x}^{j+1}-\mathbf{x}^{j}-(\boldsymbol{\Lambda}^{\mathbf{x},j}\boldsymbol{\mu}_{\text{{f}}}^{j}+\mathbf{f}^{\mathbf{x},j})\Delta t\right),$ (3.3)
$\mathbf{R}_{\text{{s}}}^{j,n}$ $=\mathbf{R}_{\text{f}}^{j}+\mathbf{E}^{j}\left(\mathbf{R}_{\text{{s}}}^{j+1,n}(\mathbf{E}^{j})^{\dagger}-(\mathbf{I}_{l\times l}+\boldsymbol{\Lambda}^{\mathbf{y},j}\Delta t)\mathbf{R}_{\text{f}}^{j}\right)-\mathbf{F}^{j}\boldsymbol{\Lambda}^{\mathbf{x},j}\mathbf{R}_{\text{f}}^{j}\Delta t,$ (3.4)
@@
where $\mathbf{I}_{l\times l}$ is the $l\times l$ identity matrix and the auxiliary matrices appearing in (3.3)–(3.4) are given up to leading-order $O(\Delta t)$, since we are using the Euler-Maruyama scheme to temporally discretize the CGNS of equations, as
$\mathbf{E}^{j}$ $:=\mathbf{I}_{l\times l}+\left[(\boldsymbol{\Sigma}^{\mathbf{y}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})^{j}\big{(}(\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})^{j}\big{)}^{-1}\mathbf{G}^{\mathbf{x},j}-\mathbf{G}^{\mathbf{y},j}\right]\Delta t+O(\Delta t^{2})\in\mathbb{C}^{l\times l},$ (3.5)
@@
$\begin{split}\mathbf{F}^{j}&:=-\mathbf{R}_{\text{{f}}}^{j}\Big{[}(\mathbf{K}^{j})^{\dagger}+\left((\mathbf{G}^{\mathbf{x},j})^{\dagger}\mathbf{K}^{j}\mathbf{R}_{\text{{f}}}^{j}(\mathbf{K}^{j})^{\dagger}-(\mathbf{R}_{\text{{f}}}^{j})^{-1}(\mathbf{H}^{j})^{\dagger}\mathbf{R}_{\text{{f}}}^{j}(\mathbf{K}^{j})^{\dagger}+(\boldsymbol{\Lambda}^{\mathbf{y},j})^{\dagger}(\mathbf{K}^{j})^{\dagger}\right)\Delta t\\ &\hskip 42.67912pt-(\boldsymbol{\Lambda}^{\mathbf{x},j})^{\dagger}\left(\big{(}(\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})^{j}\big{)}^{-1}+\mathbf{K}^{j}\mathbf{R}_{\text{{f}}}^{j}(\mathbf{K}^{j})^{\dagger}\Delta t\right)\Big{]}+O(\Delta t^{2})\in\mathbb{C}^{l\times k},\end{split}$ (3.6)
@@
$\begin{gathered}\mathbf{G}^{\mathbf{x},j}:=\boldsymbol{\Lambda}^{\mathbf{x},j}+(\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{y}})^{j}(\mathbf{R}_{\text{{f}}}^{j})^{-1}\in\mathbb{C}^{k\times l},\quad\mathbf{G}^{\mathbf{y},j}:=\boldsymbol{\Lambda}^{\mathbf{y},j}+(\boldsymbol{\Sigma}^{\mathbf{y}}\circ\boldsymbol{\Sigma}^{\mathbf{y}})^{j}(\mathbf{R}_{\text{{f}}}^{j})^{-1}\in\mathbb{C}^{l\times l},\\ \mathbf{H}^{j}:=(\mathbf{R}_{\text{{f}}})^{-1}\left(\boldsymbol{\Lambda}^{\mathbf{y},j}\mathbf{R}_{\text{f}}^{j}+\mathbf{R}_{\text{f}}^{j}(\boldsymbol{\Lambda}^{\mathbf{y},j})^{\dagger}+(\boldsymbol{\Sigma}^{\mathbf{y}}\circ\boldsymbol{\Sigma}^{\mathbf{y}})^{j}\right)\in\mathbb{C}^{l\times l},\quad\mathbf{K}^{j}:=\big{(}(\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})^{j}\big{)}^{-1}\mathbf{G}^{\mathbf{x},j}\in\mathbb{C}^{k\times l}.\end{gathered}$ (3.7)
@@
$\mathbf{E}^{j}$ $:=\mathbf{I}_{l\times l}-\mathbf{G}^{\mathbf{y},j}\Delta t+O(\Delta t^{2}),$ (3.8)
$\mathbf{F}^{j}$ $:=\mathbf{G}^{\mathbf{y},j}\mathbf{R}_{\text{f}}^{j}(\boldsymbol{\Lambda}^{\mathbf{x},j})^{\dagger}\big{(}(\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})^{j}\big{)}^{-1}\Delta t+O(\Delta t^{2}).$ (3.9)
@@
Let $(\mathbf{x}(t),\mathbf{y}(t))$ satisfy (2.1)–(2.2) and assume the validity of the assumptions in Theorem 2.3 and all regularity conditions outlined in Appendix A. Suppose now the observational data for the observed process, $\mathbf{x}^{0},\mathbf{x}^{1},\mathbf{x}^{2},\ldots$, are given sequentially. When a new observation, denoted by $\mathbf{x}^{n}$ for $n\in\mathbb{N}$, becomes available, it is utilized to update all the existing optimal state estimates at time instants $t_{j}$ for $10\leq j\leq n-$ 1 and it thens provides a new state estimate at $j=n$. The discrete smoother posterior distribution $\mathbb{P}\big{(}\mathbf{y}^{j}\big{|}\mathbf{x}^{s},0\leq s\leq n\big{)}$ is conditionally Gaussian,
$\mathbb{P}\big{(}\mathbf{y}^{j}\big{|}\mathbf{x}^{s},0\leq s\leq n\big{)}\overset{{\,\rm d}}{\sim}\mathcal{N}_{l}(\boldsymbol{\mu}_{\text{s}}^{j,n},\mathbf{R}_{\text{s}}^{j,n}),$ (3.10)
and the conditional mean $\boldsymbol{\mu}_{\text{s}}^{j,n}$ and conditional covariance $\mathbf{R}_{\text{s}}^{j,n}$ for $n\in\mathbb{N}$ and $10\leq j\leq n-$ 1 satisfy the following recursive backward difference equations:
$\boldsymbol{\mu}_{\text{s}}^{j,n}$ $=\boldsymbol{\mu}_{\text{s}}^{j,n-1}+\mathbf{D}^{j,n-2}\left(\boldsymbol{\mu}_{\text{s}}^{n-1,n}-\boldsymbol{\mu}_{\text{f}}^{n-1}\right),$ (3.11)
$\mathbf{R}_{\text{s}}^{j,n}$ $=\mathbf{R}_{\text{s}}^{j,n-1}+\mathbf{D}^{j,n-2}\left(\mathbf{R}_{\text{s}}^{n-1,n}-\mathbf{R}_{\text{f}}^{n-1}\right)(\mathbf{D}^{j,n-1})^{\dagger},$ (3.12)
@@
[…]
@@
$\begin{split}\mathbf{b}^{n-1}&:=\boldsymbol{\mu}_{\text{f}}^{n-1}-\mathbf{E}^{n-1}\left((\mathbf{I}_{l\times l}+\boldsymbol{\Lambda}^{\mathbf{y},n-1}\Delta t)\boldsymbol{\mu}_{\text{{f}}}^{n-1}+\mathbf{f}^{\mathbf{y},n-1}\Delta t\right)\\ &\hskip 42.67912pt+\mathbf{F}^{n-1}\left(\mathbf{x}^{n}-\mathbf{x}^{n-1}-(\boldsymbol{\Lambda}^{\mathbf{x},n-1}\boldsymbol{\mu}_{\text{{f}}}^{n-1}+\mathbf{f}^{\mathbf{x},n-1})\Delta t\right),\end{split}$ (3.16)
@@
Based on this theorem, the algorithm associated with the online forward-in-time discrete smoother update equations (3.11)–(3.12) is outlined in Algorithm 1 in the form of pseudocode. Notice that in Algorithm 1 we use a compact expression for the update matrix $\mathbf{D}^{j,n-2}$,

### v2

@@ 2.2 Analytically Solvable Filter and Smoother Posterior Distributions
Then, for any $T\geq t\geq 0$ ($t$ running backward), the $\mathcal{F}_{T}^{\mathbf{x}}$-measurable Gaussian statistics of the smoother posterior distribution
$\mathbb{P}\big(\mathbf{y}(t)\big|\mathbf{x}(s),s\leq T\big)\overset{{\,\rm d}}{\sim}\mathcal{N}_{l}\big(\boldsymbol{\mu}_{\text{{s}}}(t),\mathbf{R}_{\text{{s}}}(t)\big),$
defined as
$\boldsymbol{\mu}_{\text{{s}}}(t):=\mathbb{E}\big[\mathbf{y}(t)\big|\mathbf{x}(s),s\leq T\big]\quad\text{and}\quad\mathbf{R}_{\text{{s}}}(t):=\mathbb{E}\big[(\mathbf{y}(t)-\boldsymbol{\mu}_{\text{{s}}}(t))(\mathbf{y}(t)-\boldsymbol{\mu}_{\text{{s}}}(t))^{\dagger}\big|\mathbf{x}(s),s\leq T\big],$
are the unique continuous solutions to the system of optimal nonlinear smoother backward differential equations:
$\overleftarrow{{\,\rm d}\boldsymbol{\mu}_{\text{{s}}}}$ $=-\left(\boldsymbol{\Lambda}^{\mathbf{y}}\boldsymbol{\mu}_{\text{{s}}}+\mathbf{f}^{\mathbf{y}}-\mathbf{B}\mathbf{R}_{\text{{f}}}^{-1}(\boldsymbol{\mu}_{\text{{f}}}-\boldsymbol{\mu}_{\text{{s}}})\right){\,\rm d}t+(\boldsymbol{\Sigma}^{\mathbf{y}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})(\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})^{-1}\big(\overleftarrow{{\,\rm d}\mathbf{x}}+(\boldsymbol{\Lambda}^{\mathbf{x}}\boldsymbol{\mu}_{\text{{s}}}+\mathbf{f}^{\mathbf{x}}){\,\rm d}t\big),$ (2.8)
$\overleftarrow{{\,\rm d}\mathbf{R}_{\text{{s}}}}$ $=-\big((\mathbf{A}+\mathbf{B}\mathbf{R}_{\text{{f}}}^{-1})\mathbf{R}_{\text{{s}}}+\mathbf{R}_{\text{{s}}}(\mathbf{A}+\mathbf{B}\mathbf{R}_{\text{{f}}}^{-1})^{\dagger}-\mathbf{B}\big){\,\rm d}t.$ (2.9)
The backward-arrow notation in (2.8)–(2.9) is to be understood as:
$\overleftarrow{{\,\rm d}\boldsymbol{\mu}_{\text{{s}}}}:=\lim_{\Delta t\to 0}\big(\boldsymbol{\mu}_{\text{{s}}}(t)-\boldsymbol{\mu}_{\text{{s}}}(t+\Delta t)\big),\quad\overleftarrow{{\,\rm d}\mathbf{R}_{\text{{s}}}}:=\lim_{\Delta t\to 0}\big(\mathbf{R}_{\text{{s}}}(t)-\mathbf{R}_{\text{{s}}}(t+\Delta t)\big),\quad\overleftarrow{{\,\rm d}\mathbf{x}}:=\lim_{\Delta t\to 0}\big(\mathbf{x}(t)-\mathbf{x}(t+\Delta t)\big).$
In other words, the notation $\overleftarrow{\frac{{\,\rm d}\cdot}{{\,\rm d}t}}$ corresponds to the negative of the usual derivative, which means the system in (2.8)–(2.9) is to be solved backward over $[0,T]$. The “starting” values for the smoother posterior statistics, $\big(\boldsymbol{\mu}_{\text{{s}}}(T),\mathbf{R}_{\text{{s}}}(T)\big)$, are the same as those of the corresponding filter estimates at the endpoint $t=T$, $\big(\boldsymbol{\mu}_{\text{{f}}}(T),\mathbf{R}_{\text{{f}}}(T)\big)$.
@@
This is Theorem 12.10 in Liptser & Shiryaev [33]. For the analogous result in the case of discrete time, see Theorem 13.12, which outlines the corresponding optimal recursive nonlinear smoother backward difference equations, with the respective sufficient assumptions given in Subchapters 13.2.1 and 13.3.8. For a martingale-free proof see Theorem 3 in Andreou & Chen [55] (although it additionally requires assumptions (10) and (13) found in Appendix A, as to pass from the discrete-time smoother to the continuous one via a formal limit). ∎
@@ 3 Optimal Online Smoother
As aforementioned, the standard smoothing procedure involves a forward pass using a filtering method followed by a backward pass to obtain the optimal smoother state estimation [14, 15, 16, 17]. Recalculating the smoother statistics from scratch whenever new observations become available is computationally and storage-intensive. Therefore, an online version of the optimal nonlinear smoother, requiring only forward-in-time updates, is highly desirable, as it can update estimated states sequentially with new data. Fortunately, the CGNS framework provides exact, closed-form solutions for deriving an online smoother. These facilitate efficiency and the understanding of the mathematical and numerical properties of the CGNS.
@@ 3.1 Optimal Online Forward-In-Time Discrete Smoother
In what follows, the superscript notation $\cdot^{\,j}$ is used to denote the discrete approximation to the continuous form of the respective vector or matrix functional when evaluated on $t_{j}$, for example $\boldsymbol{\Lambda}^{\mathbf{x},j}:=\boldsymbol{\Lambda}^{\mathbf{x}}(t_{j},\mathbf{x}(t_{j}))$. We write the CGNS of equations (2.1)–(2.2) in a discrete fashion using the Euler–Maruyama scheme [61, 62]. Then the time discretization of (2.1)–(2.2) is simply given by
@@
$\boldsymbol{\mu}_{\text{{s}}}^{j,n}:=\mathbb{E}\big[\mathbf{y}^{j}\big|\mathbf{x}^{s},s\leq n\big],\quad\mathbf{R}_{\text{{s}}}^{j,n}:=\mathrm{Cov}\big(\mathbf{y}^{j},\mathbf{y}^{j}\big|\mathbf{x}^{s},s\leq n\big)=\mathbb{E}\big[(\mathbf{y}^{j}-\boldsymbol{\mu}_{\text{{s}}}^{j,n})(\mathbf{y}^{j}-\boldsymbol{\mu}_{\text{{s}}}^{j,n})^{\dagger}\big|\mathbf{x}^{s},s\leq n\big].$
@@
Notice how we explicitly note the dependence of the smoother state estimates on the length of the observational period. As is known by the discrete-time counterpart of Theorem 2.1 (see Theorem 13.3 in Liptser & Shiryaev [33]), the smoother posterior distribution $\mathbb{P}\big(\mathbf{y}^{j}\big|\mathbf{x}^{s},s=0,\ldots,n\big)$ is ($\mathbb{P}$-almost surely) conditionally Gaussian, $\mathcal{N}_{l}\big(\boldsymbol{\mu}_{\text{{s}}}^{j,n},\mathbf{R}_{\text{{s}}}^{j,n}\big)$, and it is possible to show that the conditional mean $\boldsymbol{\mu}_{\text{{s}}}^{j,n}$ and conditional covariance $\mathbf{R}_{\text{{s}}}^{j,n}$ of the discrete smoother at time step $t_{j}$ when conditioning up to the $n$-th observation satisfy the following recursive backward difference equations for $n\in\mathbb{N}$ and $j=0,1,\ldots,n-1$ (under the regularity conditions (1)(13) in Appendix A and all other assumptions thus far) [55]:
$\boldsymbol{\mu}_{\text{{s}}}^{j,n}$ $=\boldsymbol{\mu}_{\text{{f}}}^{j}+\mathbf{E}^{j}\big(\boldsymbol{\mu}_{\text{{s}}}^{j+1,n}-(\mathbf{I}_{l\times l}+\boldsymbol{\Lambda}^{\mathbf{y},j}\Delta t)\boldsymbol{\mu}_{\text{{f}}}^{j}-\mathbf{f}^{\mathbf{y},j}\Delta t\big)+\mathbf{F}^{j}\big(\mathbf{x}^{j+1}-\mathbf{x}^{j}-(\boldsymbol{\Lambda}^{\mathbf{x},j}\boldsymbol{\mu}_{\text{{f}}}^{j}+\mathbf{f}^{\mathbf{x},j})\Delta t\big),$ (3.3)
$\mathbf{R}_{\text{{s}}}^{j,n}$ $=\mathbf{R}_{\text{f}}^{j}+\mathbf{E}^{j}\big(\mathbf{R}_{\text{{s}}}^{j+1,n}(\mathbf{E}^{j})^{\dagger}-(\mathbf{I}_{l\times l}+\boldsymbol{\Lambda}^{\mathbf{y},j}\Delta t)\mathbf{R}_{\text{f}}^{j}\big)-\mathbf{F}^{j}\boldsymbol{\Lambda}^{\mathbf{x},j}\mathbf{R}_{\text{f}}^{j}\Delta t,$ (3.4)
@@
where $\mathbf{I}_{l\times l}$ is the $l\times l$ identity matrix and the auxiliary matrices appearing in (3.3)–(3.4) are given up to leading-order $O(\Delta t)$, since we are using the Euler–Maruyama scheme to temporally discretize the CGNS of equations, as
$\mathbf{E}^{j}$ $:=\mathbf{I}_{l\times l}+\big((\boldsymbol{\Sigma}^{\mathbf{y}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})^{j}\big((\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})^{j}\big)^{-1}\mathbf{G}^{\mathbf{x},j}-\mathbf{G}^{\mathbf{y},j}\big)\Delta t+O(\Delta t^{2})\in\mathbb{C}^{l\times l},$ (3.5)
@@
$\begin{split}\mathbf{F}^{j}&:=-\mathbf{R}_{\text{{f}}}^{j}\Big((\mathbf{K}^{j})^{\dagger}+\big((\mathbf{G}^{\mathbf{x},j})^{\dagger}\mathbf{K}^{j}\mathbf{R}_{\text{{f}}}^{j}(\mathbf{K}^{j})^{\dagger}-(\mathbf{R}_{\text{{f}}}^{j})^{-1}(\mathbf{H}^{j})^{\dagger}\mathbf{R}_{\text{{f}}}^{j}(\mathbf{K}^{j})^{\dagger}+(\boldsymbol{\Lambda}^{\mathbf{y},j})^{\dagger}(\mathbf{K}^{j})^{\dagger}\big)\Delta t\\ &\hskip 42.67912pt-(\boldsymbol{\Lambda}^{\mathbf{x},j})^{\dagger}\big(\big((\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})^{j}\big)^{-1}+\mathbf{K}^{j}\mathbf{R}_{\text{{f}}}^{j}(\mathbf{K}^{j})^{\dagger}\Delta t\big)\Big)+O(\Delta t^{2})\in\mathbb{C}^{l\times k},\end{split}$ (3.6)
@@
$\begin{gathered}\mathbf{G}^{\mathbf{x},j}:=\boldsymbol{\Lambda}^{\mathbf{x},j}+(\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{y}})^{j}(\mathbf{R}_{\text{{f}}}^{j})^{-1}\in\mathbb{C}^{k\times l},\quad\mathbf{G}^{\mathbf{y},j}:=\boldsymbol{\Lambda}^{\mathbf{y},j}+(\boldsymbol{\Sigma}^{\mathbf{y}}\circ\boldsymbol{\Sigma}^{\mathbf{y}})^{j}(\mathbf{R}_{\text{{f}}}^{j})^{-1}\in\mathbb{C}^{l\times l},\\ \mathbf{H}^{j}:=(\mathbf{R}_{\text{{f}}})^{-1}\big(\boldsymbol{\Lambda}^{\mathbf{y},j}\mathbf{R}_{\text{f}}^{j}+\mathbf{R}_{\text{f}}^{j}(\boldsymbol{\Lambda}^{\mathbf{y},j})^{\dagger}+(\boldsymbol{\Sigma}^{\mathbf{y}}\circ\boldsymbol{\Sigma}^{\mathbf{y}})^{j}\big)\in\mathbb{C}^{l\times l},\quad\mathbf{K}^{j}:=\big((\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})^{j}\big)^{-1}\mathbf{G}^{\mathbf{x},j}\in\mathbb{C}^{k\times l}.\end{gathered}$ (3.7)
@@
$\mathbf{E}^{j}:=\mathbf{I}_{l\times l}-\mathbf{G}^{\mathbf{y},j}\Delta t+O(\Delta t^{2}),\quad\mathbf{F}^{j}:=\mathbf{G}^{\mathbf{y},j}\mathbf{R}_{\text{f}}^{j}(\boldsymbol{\Lambda}^{\mathbf{x},j})^{\dagger}\big((\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})^{j}\big)^{-1}\Delta t+O(\Delta t^{2}).$ (3.8)
@@
Let $\big(\mathbf{x}(t),\mathbf{y}(t)\big)$ satisfy (2.1)–(2.2) and assume the validity of the assumptions in Theorem 2.3 and all regularity conditions outlined in Appendix A, (1)(13). Suppose now the observational data for the observed process, $\mathbf{x}^{0},\mathbf{x}^{1},\mathbf{x}^{2},\ldots$, are given sequentially. When a new observation, denoted by $\mathbf{x}^{n}$ for $n\in\mathbb{N}$, becomes available, it is utilized to update all the existing optimal state estimates at time instants $t_{j}$ for $0\leq j\leq n-1$ and it then provides a new state estimate at $j=n$. The discrete smoother posterior distribution $\mathbb{P}\big(\mathbf{y}^{j}\big|\mathbf{x}^{s},0\leq s\leq n\big)$ is conditionally Gaussian,
$\mathbb{P}\big(\mathbf{y}^{j}\big|\mathbf{x}^{s},0\leq s\leq n\big)\overset{{\,\rm d}}{\sim}\mathcal{N}_{l}\big(\boldsymbol{\mu}_{\text{s}}^{j,n},\mathbf{R}_{\text{s}}^{j,n}\big),$ (3.9)
and the conditional mean $\boldsymbol{\mu}_{\text{s}}^{j,n}$ and conditional covariance $\mathbf{R}_{\text{s}}^{j,n}$ for $n\in\mathbb{N}$ and $0\leq j\leq n-1$ satisfy the following recursive backward difference equations:
$\boldsymbol{\mu}_{\text{s}}^{j,n}$ $=\boldsymbol{\mu}_{\text{s}}^{j,n-1}+\mathbf{D}^{j,n-2}\big(\boldsymbol{\mu}_{\text{s}}^{n-1,n}-\boldsymbol{\mu}_{\text{f}}^{n-1}\big),$ (3.10)
$\mathbf{R}_{\text{s}}^{j,n}$ $=\mathbf{R}_{\text{s}}^{j,n-1}+\mathbf{D}^{j,n-2}\big(\mathbf{R}_{\text{s}}^{n-1,n}-\mathbf{R}_{\text{f}}^{n-1}\big)(\mathbf{D}^{j,n-2})^{\dagger},$ (3.11)
@@
$\mathbf{D}^{n,n-1}$ $:=\mathbf{I}_{l\times l},\hskip 28.45274pt$ $\text{ for }n\in\mathbb{N},$
$\mathbf{D}^{n-1,n-1}$ $:=\mathbf{E}^{n-1},\hskip 28.45274pt$ $\text{ for }n\in\mathbb{N},$ (3.12)
$\mathbf{D}^{j,n-1}$ $:=\mathbf{D}^{j,n-2}\mathbf{E}^{n-1},\hskip 28.45274pt$ $\text{ for }n\in\mathbb{N}_{\geq 2},\text{ and }j=0,1,\ldots,n-2,$
@@
$\begin{split}\mathbf{b}^{n-1}&:=\boldsymbol{\mu}_{\text{f}}^{n-1}-\mathbf{E}^{n-1}\big((\mathbf{I}_{l\times l}+\boldsymbol{\Lambda}^{\mathbf{y},n-1}\Delta t)\boldsymbol{\mu}_{\text{{f}}}^{n-1}+\mathbf{f}^{\mathbf{y},n-1}\Delta t\big)\\ &\hskip 42.67912pt+\mathbf{F}^{n-1}\big(\mathbf{x}^{n}-\mathbf{x}^{n-1}-(\boldsymbol{\Lambda}^{\mathbf{x},n-1}\boldsymbol{\mu}_{\text{{f}}}^{n-1}+\mathbf{f}^{\mathbf{x},n-1})\Delta t\big),\end{split}$ (3.15)
@@
Based on this theorem, the algorithm associated with the online forward-in-time discrete smoother update equations (3.10)–(3.11) is outlined in Algorithm 1. There, we use a compact expression for the update matrix $\mathbf{D}^{j,n-2}$,